Fermi National Laboratory
 

Accelerator Update for November 24, 2006 - November 29, 2006

 

Friday November 24

The day shift began with Operations monitoring store 5089 (115.5E30) and a Recycler stash (135.9E10), with the Antiproton Source (Pbar) stacking (68.8E10), with Linac, Booster, the Recycler, and the Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

At 10:36 PM, cryo system technicians at CHL reported that they had lost a cold box. CDF and D0 put their detectors in standby.

The Recycler completed the last of six stashes of antiprotons for the day at 11 PM.

The Cryo Coordinator reported at 11:21 PM that CHL had recovered.

Operations terminated store 5089 at 11:29 PM.

At 11:42 PM, Operations reported problems with the Tevatron (TeV) collimators at D4 and E0. Operators investigated and found a VME with no power.

Saturday November 25

The midnight shift began with Operations monitoring a stash, with Pbar stacking, with cryo techs investigating a cold box problem, with operators making a TeV VME power supply repair, and with MiniBooNE and NuMI taking beam.

CHL techs reported at 12:19 AM that they found a failed cold box temperature transmitter.

Operations reported at 2:47 AM that they replaced the VME power supply for the collimators.

CHL techs reported at 3:15 AM that they had replaced the failed temperature transmitter.

Operations established a 6x0 (6 bunches of protons and 0 bunches of antiprotons) study-store in the TeV at 4:43 AM.

Operations put the TeV into shot setup 5:08 AM.

Operations established store 5092 at 7:04 AM with an initial luminosity of 221.04E30.

The Recycler had not been stashing due to an instrumentation problem. Techs reported at 9:23 AM that they would have to access MI to make repairs. Operators prepared MI for access at 9:59 AM.

Instrumentation techs completed their repairs and were out of the tunnel by 10:39 AM.

The Recycler resumed stashing at 12:27 PM.

The Recycler completed its last of six stashes for Saturday at 11:38 PM.

Sunday November 26

The midnight shift began with Operations monitoring store 5092 and a stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations terminated store 5092 at 7:58 AM.

Operations put the TeV into shot setup at 8:17 AM.

At 10:07 AM, Operations reported that both beams in the TeV had disappeared before a store was established. There was no quench or abort.

Operations investigated and reported at 11:16 AM that a TeV vertical dipole power supply (VC25) had tripped and caused the beam to drop out. Operators had the power supply replaced by 11:53 AM.

At 12:04 PM, the TeV sector A1 cryo system had a Dewar problem and the system was unstable. Experts investigated and found a vacuum problem. Vacuum techs replaced a vacuum pump and installed two portable pumps on the A1 cryo system. The system was ready to go by 3:36 PM.

Operations put the TeV through a wet squeeze at 5:45 PM.

Operations put the TeV into shot setup at 7:29 PM.

Operations established store 5095 at 9:19 PM with an initial luminosity of 113.8E30.

The Recycler completed the last of seven stashes for Sunday at 10:25 PM.

Monday November 27

The midnight shift began with Operations monitoring store 5095 and a stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 8:17 AM, an expert reported finding a small LCW leak on Debuncher Lambertson power supply (D:ELAM). This leak will get repaired later.

At 9:52 AM, Operations found an LCW leak on a MI-40 quadrupole power supply.

An expert repaired the leak on D:ELAM at 11:48 AM.

Operations terminated store 5095 at 2:38 PM.

Operations established store 5096 at 5:05 PM with an initial luminosity of 191.97E30.

The Recycler completed the last and seventh stash for Monday at 11:37 PM.

Tuesday November 28

The midnight shift began with Operations monitoring store 5096 and a stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 10:45 AM, Operations received a fire alarm for MI-20.

At 10:56 AM, Operations turned off MI. Smoke was seen in the stairwell.

At 11:14 AM, Operations pulled a key so that the Fire Department could enter the tunnel. This caused the loss of the Recycler stash.

At 11:28 AM, the Fire Department reported that a lighting ballast in the stairwell had failed. Smoke never made it into the MI tunnel. The Duty Electrician was called in to repair the ballast. The repair was completed by 1 PM.

Operations had beam to MiniBooNE and Pbar by 1:21 PM.

The Recycler completed the last of eight stashes for Tuesday at 11:56 PM.

Wednesday November 29

The midnight shift began with Operations monitoring store 5096 and a stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

The Recycler completed the fourth and last stash for the midnight shift at 7:53 AM.

Operations terminated store 5096 at 8:04 AM.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Pbar and MI experts to conduct studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/29/2006    email Fermilab